For a deleterious mutation with s = −0.001, what is the probability of fixation in a population of size (a) Ne = 10, 000, (b) Ne = 1, 000, and (c) Ne = 100.
Answer #1

Answer: Probability of fixation (P) = 2 Sooool Formula = 1 - 0-25 1- -4 NS (a) N = 10,000 P= 1 e 240.00l - 4 x 10,000 xooool(e) N=100 - 240.001 p - l- én I-e-44100 X 0.001 P = 6.06410-3